OpenStudy (anonymous):

how do you solve a=c/2b for c

OpenStudy (anonymous):

2ab=c

OpenStudy (anonymous):

multiply both sides by 2b since its divided

OpenStudy (accessdenied):

In general, when you want to solve for another variable, you focus on that variable: \( a = \dfrac{\color{goldenrod}{c}}{2b} \) Solve for c. To isolate the item we want to solve for, we apply an inverse operation that cancels what is already there. Addition and Subtraction of the same quantity cancel Division and Multiplication of the same quantity cancel Note that 2b is divided off of the c. So we multiply both sides by the 2b, the inverse operation. \( a \times 2b = \dfrac{\color{goldenrod}c}{\cancel{2b}} \times \cancel{2b} \) \( a \times 2b = \color{goldenrod}c \) Now we have the value of c in terms of the other variables! We can rearrange this by commutative property to the more pleasing form, \( 2ab = c \).